# FrameOracle: Learning What to See and How Much to See in Videos

### v1

@@ Abstract
Vision-language models (VLMs) have advanced video understanding, but their performance is limited by the number of input frames they can process. Existing frame sampling strategies, such as uniform or fixed-budget selection, often fail to adapt to variations in information density or task complexity, resulting in inefficiency and information loss. To address this, we present FrameOracle, a lightweight and plug-and-play module that predicts both (1) which frames are most relevant to a given query and (2) how many frames are needed. FrameOracle is trained using a four-stage curriculum, with the first three stages relying on weak proxy signals such as cross-modal similarity. In the final stage, it leverages stronger supervision from a new dataset we introduce, FrameOracle-41K, the first large-scale VideoQA collection to provide keyframe annotations specifying the minimal set of frames required to answer each question. Extensive experiments across five VLMs and six benchmarks demonstrate that FrameOracle reduces 16-frame inputs to an average of 10.4 frames without any loss in accuracy. When starting from 64-frame candidates, it reduces the input to an average of 13.9 frames while improving accuracy by 1.4%, achieving state-of-the-art efficiency-accuracy trade-offs for scalable video understanding.
@@ 1 Introduction
Rapid advances in large language models (LLMs) (Stiennon et al., 2020; Gao et al., 2022; Yang et al., 2024) have enabled vision-language models (VLMs) to integrate visual understanding with strong linguistic reasoning (Zhang et al., 2024; Bai et al., 2025; Zhang et al., 2025a). This makes VLMs highly effective for complex video tasks such as question answering (Zhang et al., 2023; Lin et al., 2024a; b; Zhao et al., 2024; Xiao et al., 2025), summarization (Hua et al., 2025; Lee et al., 2025), and instruction following (Ren et al., 2024; Qian et al., 2024). A key challenge, however, is the large volume of data these models must process. Processing every video frame is computationally expensive, making efficient frame sampling essential (Hu et al., 2025). Most VLMs currently rely on simple approaches, such as uniform sampling at a fixed frame rate or selecting a fixed number of frames. While easy to implement, these methods have clear drawbacks: in long videos, they may miss crucial information, whereas in short videos, they often introduce redundant frames that waste resources, distract the model, and obscure key moments.
To mitigate this, a growing body of work has explored keyframe selection methods (Liu et al., 2025; Park et al., 2024; Tang et al., 2025; Zhang et al., 2025d). These approaches aim to identify a subset of frames that preserves semantic content while reducing redundancy. However, most existing methods assume a fixed, preset number of keyframes, ignoring the fact that the optimal number of frames varies across videos and queries. For example, short action-centric questions (e.g., whether a ball crosses a line in sports footage) may be resolved with just a handful of frames, while long-form narrative reasoning (e.g., inferring character intentions in a film) often requires a substantially larger set of frames. A few recent methods enable adaptive frame selection, but their adaptivity remains limited. In some cases, the selector is trained jointly with the backbone VLM (Buch et al., 2025), making it non-transferable to other models. In others, adaptivity is achieved via threshold-based filtering at inference, retaining only keyframes above a preset reward threshold. While this produces variable frame counts, it is not explicitly optimized during training, reducing effectiveness and generalizability. This raises a fundamental question: How can we design a selector that identifies the most relevant frames for a given query and determines how many are needed, while generalizing across different VLMs?
@@
To this end, we propose FrameOracle, a lightweight, plug-and-play frame selector that can be integrated with arbitrary VLMs. Unlike prior approaches that fix the number of frames in advance or require co-training with a specific backbone, FrameOracle jointly predicts (1) the importance of each frame relative to the query and (2) the number of frames to retain. The module is trained with a four-stage curriculum. The first three stages rely on weak proxy signals, such as cross-modal similarity and leave-one-out loss degradation. The final stage leverages stronger supervision from a new dataset we create, FrameOracle-41K, a large-scale VideoQA dataset with 40,992 examples and the first to provide keyframe annotations specifying the minimal frames required to answer each question. Unlike tasks such as object detection (Lin et al., 2014) or captioning (Xiong et al., 2024), no existing dataset provides ground-truth annotations identifying the keyframes. FrameOracle dynamically adapts its selections based on both video content and the prompt, operating seamlessly as a pre-processing module for any downstream VLM.
@@
  • We propose FrameOracle, a lightweight and plug-and-play frame selector that dynamically predicts both which frames are most relevant and how many are needed.
  • To facilitate training, we introduce FrameOracle-41K, the first large-scale VideoQA dataset with keyframe annotations, specifying the minimal set of frames needed to answer each question.
  • We conduct extensive experiments across five VLMs and six benchmarks, showing that FrameOracle reduces 16-frame inputs to an average of 10.4 frames without any loss in accuracy. When starting from 64-frame candidates, it reduces the input to an average of 13.9 frames while improving accuracy by 1.4%, achieving state-of-the-art efficiency-accuracy trade-offs for scalable video understanding.
@@ 2 Related Work
Keyframe Selection for Video Understanding. Most existing keyframe selection methods assume a fixed frame budget: they rank candidate frames by visual–linguistic relevance or temporal salience and then retain the top-$k$ subset (Liang et al., 2024; Tan et al., 2024; Yu et al., 2025; Liu et al., 2025; Fang et al., 2025; Tang et al., 2025). Beyond this fixed-budget paradigm, some work has explored adaptive frame selection. These approaches fall into two categories. The first are agent-based methods, where large multimodal models act as decision-makers that iteratively analyze videos. For instance, VCA (Yang et al., 2025) combines curiosity-driven exploration with tree search to identify informative segments, while AKeyS (Fan et al., 2025) leverages a language agent to heuristically expand video segments and decide both which frames to retain and when to stop. However, such methods are computationally expensive due to repeated agent calls. The second category comprises approaches that require co-training with a specific VLM backbone (Buch et al., 2025; Yu et al., 2023; Guo et al., 2025), which restricts their portability. In contrast, FrameOracle is adaptive, lightweight, and model-agnostic: it learns to jointly predict which frames are relevant and how many to retain, while remaining plug-and-play across diverse VLMs.
@@
3 FrameOracle-41K Dataset
We introduce FrameOracle-41K, the first VideoQA dataset that provides keyframe annotations, specifying the minimal set of frames needed to answer each question. The corpus contains 40,992 video–question pairs spanning diverse scenes and durations. In contrast to existing VideoQA datasets, which provide only ground-truth answers and, in some cases, coarse temporal spans in the video, FrameOracle-41K records, for each instance, the minimal number of frames needed to answer the question with high confidence, along with the keyframes that constitute the necessary evidence. Below, we describe our data collection and generation pipeline, the verification and filtering procedures used to retain high-quality data, and the key statistics of the resulting dataset.
@@ 3.1 Data Gathering and Processing
All video–question pairs in FrameOracle-41K are sourced from LLaVA-Video-178K (Zhang et al., 2024), a large-scale VideoQA dataset that covers a wide range of scenarios and activities. From this corpus, we first select nearly 100k videos, each 2–3 minutes long, balancing adequate temporal context with a manageable annotation effort. We then apply a two-stage process to create the final dataset. Stage I (agent-based keyframe mining) iteratively explores each video using a multimodal agent. At the end of exploration, the agent outputs a predicted answer with its confidence and assigns a relevance score to every visited frame. Stage II (filtering and verification) first removes frames with relevance scores below a threshold, then verifies sufficiency by requiring three independent VLMs to answer correctly using only the remaining keyframes. Only instances that pass this verification are retained, finalizing the minimal sufficient frame count. An overview of the pipeline is shown in Figure 1. Appendix B provides an example of a data instance in JSON format.
@@
Stage I: Agent-based Keyframe Mining. Starting from a uniformly sampled set of 64 frames, we employ an agent built on Qwen2.5-VL-72B API (Bai et al., 2025) to iteratively explore the video with respect to the given question. In the first iteration, the agent inspects three anchor frames (indices 0, 31, and 63), assigns relevance scores, and attempts an answer with a confidence estimate. It then compares the pairwise summed relevance of adjacent anchors (0+31 vs. 31+63) to decide which segment to explore next. Within the selected segment, a denser set of four anchor frames is sampled, another answer with confidence is attempted, and the same pairwise relevance comparison guides subsequent iterations. This iterative score–refine cycle continues until either the agent becomes confident enough to provide a stable answer or all frames have been examined. By the end of Stage I, the agent returns (1) its predicted answer and confidence, and (2) the complete set of frames it has inspected, each annotated with a relevance score. Any video-question pair for which the agent’s predicted answer does not match the ground-truth answer is then discarded. This mined trajectory captures both the localization of question-specific evidence and fine-grained frame-level importance signals, forming the raw candidates for the next stage.
@@
Stage II: Filtering and Verification. After obtaining candidate keyframes from Stage I, we first remove all frames with relevance scores below a threshold $\lambda$, leaving only those with stronger relevance. For each video–question pair, we then test whether the selected keyframes alone are sufficient to answer the question. Specifically, the keyframe set and the question are fed into three independent VLMs (i.e., Qwen2.5-VL-72B (Bai et al., 2025), LLaVA-OneVision-72B (Li et al., 2025), and LLaVA-Video-72B (Zhang et al., 2024)), and their predictions are compared against the ground-truth answer. Only instances for which all three models succeed using only the keyframes are retained; otherwise, the instance is discarded. This cross-model verification ensures that the released dataset contains consistent, question-grounded keyframe annotations.
(a) Distribution of video durations in FrameOracle-41K.
(b) Distribution of the number of selected keyframes per video–question pair.
Figure 2: FrameOracle-41K dataset statistics. (a) Video durations range from 2 to 3 minutes, offering enough temporal context for analysis without including unnecessary frames. (b) Keyframe annotations are sparse: most videos require fewer than 10 frames, yet the dataset also includes harder edge cases to capture more challenging scenarios.
@@ 3.2 Dataset Statistics
Through the two-stage pipeline described above, we obtain 40,992 validated video–question pairs, each annotated with keyframes and minimal sufficient frame counts. This collection forms the FrameOracle-41K dataset. Figure 2(a) shows that most videos are 2 to 3 minutes long, long enough to provide meaningful temporal context without the inefficiency and redundancy of very long videos. Figure 2(b) shows the distribution of keyframes per video–question pair. The median is about five frames, the mean is around seven, and more than 80% of samples require no more than 10 frames. A small fraction of more challenging cases, however, require 30 or more frames. These statistics highlight the challenge and value of FrameOracle-41K: models must learn to identify minimal yet sufficient evidence across diverse temporal contexts, rather than relying on dense frame sampling.
4 Method
We introduce FrameOracle, a lightweight and adaptive frame selector that dynamically determines the appropriate number of keyframes from a video, conditioned on the user prompt. FrameOracle enables efficient video understanding by providing the downstream VLM with a compact yet highly relevant subset of frames.
Since directly processing all frames of a video, $V$, is computationally expensive, we first apply uniform temporal sampling to extract a candidate set of $N$ frames, denoted as $V_{C}=\{f_{1},\dots,f_{N}\}$. This pre-sampling step acts as a coarse filter, reducing the input to a manageable size for FrameOracle (e.g., $N=64$ or $N=16$). Our goal is to learn FrameOracle, a selection policy, $\Pi_{\theta}$, parameterized by $\theta$, that operates on the candidate set. Given a candidate set $V_{C}$ and a text prompt $P$, FrameOracle selects a compact subset of frames $V_{S}\subset V_{C}$. Unlike approaches that fix the number of selected frames in advance, FrameOracle dynamically determines the subset size, $K=|V_{S}|$, as part of the selection process. Formally, $\Pi_{\theta}$ maps the pair $(V_{C},P)$ to the selected subset $V_{S}$, which is then passed to a downstream VLM, $\mathcal{M}$, to perform a reasoning task, producing an output $A=\mathcal{M}(V_{S},P)$. The objective is to train $\Pi_{\theta}$ to choose subsets that maximize the performance of $\mathcal{M}$ while keeping $K$ as small as possible.
@@ 4 Method
4.1 FrameOracle
FrameOracle, $\Pi_{\theta}$, is a neural module that learns to jointly predict frame importance and the number of frames to select from the candidate set $V_{C}$. We begin by extracting features from both the video frames and the text prompt, which serve as inputs to FrameOracle. For the candidate frame set $V_{C}$, we use a visual encoder to generate a sequence of $N$ frame embeddings. The text prompt $P$ is encoded using the native tokenizer of the downstream VLM. The FrameOracle architecture, shown in Figure 3, is composed of two main components: (1) a cross-modal fusion encoder and (2) dual prediction heads.
(1) Cross-Modal Fusion. To model the relationship between the text query and the video, we fuse the two modalities. Frame and text embeddings are first projected into a shared latent space using linear layers, and then processed by a stack of Transformer encoder layers. This architecture captures the complex interactions between the textual prompt and the visual content of the candidate frames.
(2) Dual Prediction Heads. The output of the fusion encoder is passed to two specialized heads, which form the core of our selection policy:
  • Rank Head: This head evaluates the relevance of each candidate frame to the prompt. It processes the fused feature sequence to output a scalar importance score, $s_{i}$, for each frame $f_{i}\in V_{C}$, resulting in a score vector $S=\{s_{1},\dots,s_{N}\}$.
  • K Head: This head predicts how many frames to select from the candidate set. It takes the globally aggregated features from the fusion encoder and outputs a probability distribution over a discrete set of possible values for $K$, where $K\leq N$.

### v2

@@ Abstract
Vision-language models (VLMs) advance video understanding but operate under tight computational budgets, making performance dependent on selecting a small, high-quality subset of frames. Existing frame sampling strategies, such as uniform or fixed-budget selection, fail to adapt to variations in content density or task complexity. To address this, we present FrameOracle, a lightweight, plug-and-play module that predicts both (1) which frames are most relevant to a given query and (2) how many frames are needed. FrameOracle is trained via a curriculum that progresses from weak proxy signals, such as cross-modal similarity, to stronger supervision with FrameOracle-41K, the first large-scale VideoQA dataset with validated keyframe annotations specifying minimal sufficient frames per question. Extensive experiments across five VLMs and six benchmarks show that FrameOracle reduces 16-frame inputs to an average of 10.4 frames without accuracy loss. When starting from 64-frame candidates, it reduces inputs to 13.9 frames on average while improving accuracy by 1.5%, achieving state-of-the-art efficiency–accuracy trade-offs for scalable video understanding.
Machine Learning, ICML
@@ 1 Introduction
Rapid advances in large language models (LLMs) (Stiennon et al., 2020; Gao et al., 2022; Yang et al., 2024) have enabled vision-language models (VLMs) to combine visual understanding with strong linguistic reasoning (Zhang et al., 2024; Bai et al., 2025; Zhang et al., 2025a). As a result, VLMs are highly effective for complex video tasks such as question answering (Zhang et al., 2023; Lin et al., 2024a, b; Zhao et al., 2024; Xiao et al., 2025), summarization (Hua et al., 2025; Lee et al., 2025b), and instruction following (Ren et al., 2024; Qian et al., 2024). In practice, however, these models must operate under tight computational budgets, limiting the number of video frames they can process. Under such constraints, performance depends on selecting a small, high-quality subset of frames. Most existing VLMs rely on simple strategies, such as uniform sampling at a fixed frame rate or selecting a fixed number of frames (keyframes). While easy to implement, these approaches fail to adapt to variations in content density or task complexity, often missing salient moments in long videos while introducing redundant and distracting frames in shorter ones.
To mitigate this issue, a growing body of work explores keyframe selection methods (Liu et al., 2025; Park et al., 2024; Tang et al., 2025; Zhang et al., 2025d), which aim to preserve semantic content while reducing redundancy. However, most existing approaches assume a fixed number of keyframes, overlooking the fact that the optimal frame budget varies across videos and queries. A few recent methods introduce adaptive frame selection, but their adaptivity remains limited. Some require joint training with a specific VLM backbone (Buch et al., 2025), restricting transferability, while others use heuristics at inference rather than learning frame selection end-to-end (Yu et al., 2025). These limitations raise the question: Can frame selection be designed as a plug-and-play learnable module that decides which frames to select (“what to see”) and how many (“how much to see”) for a given query?
@@
To answer this question, we introduce FrameOracle, a lightweight, backbone-agnostic frame selector that can be integrated with any VLM. Unlike prior methods that fix the number of frames or require co-training with a specific backbone, FrameOracle jointly predicts (1) the importance of each frame relative to the query and (2) how many frames to retain. It is trained progressively, starting with weak proxy signals and refined with stronger supervision from our new dataset, FrameOracle-41K, the first large-scale VideoQA dataset with keyframe annotations specifying the minimal frames required to answer each question. To our knowledge, no dataset provides ground-truth keyframes for VideoQA. FrameOracle adapts its selections to both video content and query, functioning seamlessly as a pre-processing module for downstream VLMs.
@@
  • We introduce FrameOracle, a plug-and-play frame selector that predicts which frames to see and how many frames to retain for each video and query.
  • We present FrameOracle-41K, the first large-scale VideoQA dataset with keyframe annotations specifying the minimal frames required to answer each question.
  • We conduct extensive experiments across five VLMs and six benchmarks, showing that FrameOracle reduces 16-frame inputs to 10.4 frames on average without accuracy loss, and reduces 64-frame inputs to 13.9 frames on average while improving accuracy by 1.5%, achieving state-of-the-art efficiency–accuracy trade-offs.
@@ 2 Related Work
Keyframe Selection for Video Understanding. Most existing keyframe selection methods assume a fixed frame budget: they rank candidate frames by visual–linguistic relevance or temporal salience and then retain the top-$k$ subset, a paradigm adopted by methods such as KeyVideoLLM (Liang et al., 2024), BOLT (Liu et al., 2025), MDP3 (Sun et al., 2025), and ReFoCUS (Lee et al., 2025a). Beyond this fixed-budget paradigm, adaptive frame selection has been explored in both classical video recognition and recent VLM-based settings. Early work, such as AdaFrame (Wu et al., 2019), proposes a learned policy to dynamically decide how many frames to process for video classification. More recently, adaptive selection has been revisited in the context of video-language reasoning. These approaches fall into two categories. The first are agent-based methods, where large multimodal models act as decision-makers that iteratively analyze videos. For instance, VCA (Yang et al., 2025) combines curiosity-driven exploration with tree search to identify informative segments, while AKeyS (Fan et al., 2025) leverages a language agent to heuristically expand video segments and decide both which frames to retain and when to stop. However, such methods are computationally expensive due to repeated agent calls. The second category comprises approaches that require co-training with a specific VLM backbone (Buch et al., 2025; Yu et al., 2023; Guo et al., 2025), which restricts their portability. In contrast, FrameOracle is adaptive, lightweight, and model-agnostic: it learns to jointly predict which frames are relevant and how many to retain under a cross-modal VideoQA setting, while remaining plug-and-play across diverse VLMs.
@@ 3 FrameOracle-41K Dataset
We introduce FrameOracle-41K, the first VideoQA dataset with keyframe annotations, specifying the minimal set of frames required to answer each question. The corpus contains 40,992 video-question pairs covering diverse scenes and durations. In contrast to existing VideoQA datasets, which provide only ground-truth answers and, in some cases, approximate temporal regions of the video containing the answer, FrameOracle-41K records, for each instance, the minimal number of frames needed to answer the question, along with the keyframes that provide the necessary evidence. Below, we describe our data generation pipeline, the verification and filtering procedures used to retain high-quality data, and the key statistics of the dataset.
@@ 3.1 Data Gathering and Processing
All video-question pairs in FrameOracle-41K are sourced from LLaVA-Video-178K (Zhang et al., 2024), a large-scale dataset covering a wide range of scenarios and activities. From this corpus, we first select nearly 100K videos, each 2–3 minutes long, balancing sufficient temporal context with manageable annotation effort. The final dataset is created through a two-stage process (Figure 1). Stage I (agent-based keyframe mining) automatically extracts candidate keyframes using a multimodal agent that iteratively explores each video and assigns frame-level relevance scores. Stage II (filtering and verification) first filters candidate keyframes by discarding frames with low relevance scores, and then retains only those video–question pairs for which three independent VLMs can correctly answer the question using the remaining keyframes. This ensures that the remaining keyframes contain the necessary evidence to answer each question. We further conduct a human verification on 4,000 randomly sampled instances, achieving an inter-annotator agreement of 94% and a verified accuracy of 93.3%. This confirms the reliability of the automatically generated annotations. Prompts for data generation are in Appendix C. An example dataset entry, human verification details, and dataset visualizations are shown in Appendices B, E, and H, respectively.
@@
Stage I: Agent-based Keyframe Mining. Starting from a uniformly sampled set of 64 frames, we employ an agent built on Qwen2.5-VL-72B API (Bai et al., 2025) to iteratively explore the video conditioned on the given question. In the first iteration, the agent evaluates three anchor frames (indices 0, 31, and 63), assigns each a relevance score, and attempts to answer the question with an associated confidence estimate. It then compares the summed relevance scores of the two adjacent anchor pairs ($0+31$ vs. $31+63$) and selects the higher-scoring temporal segment for further exploration. In subsequent iterations, the agent uniformly samples four new anchor frames within the selected segment, repeats relevance scoring and answer prediction, and again selects the sub-segment with the highest aggregated relevance. This coarse-to-fine procedure iteratively narrows the temporal window of interest. The process terminates when the agent reaches a confident answer or when all 64 candidate frames have been evaluated. At the end of Stage I, the agent outputs the set of visited frames with their relevance scores and discards video–question pairs whose predicted answers do not match the ground truth.
@@
Stage II: Filtering and Verification. After obtaining candidate keyframes from Stage I, we first discard frames with relevance scores below a threshold $\lambda$, leaving only those with stronger relevance. For each video–question pair, we then test whether the selected keyframes alone are sufficient to answer the question. Specifically, the keyframe set and the question are fed into three independent VLMs (i.e., Qwen2.5-VL-72B (Bai et al., 2025), LLaVA-OneVision-72B (Li et al., 2025), and LLaVA-Video-72B (Zhang et al., 2024)), and their predictions are compared with the ground-truth answer. Only instances for which all three models succeed using only the keyframes are retained. This cross-model verification ensures that the released dataset contains consistent, question-grounded keyframe annotations.
Figure 2: FrameOracle-41K question-level statistics. Left: Distribution of 16 question types across the dataset. Right: Per-type distributions of minimal sufficient keyframes.
@@ 3.2 Dataset Statistics
Our two-stage pipeline produces 40,992 video–question pairs, forming the FrameOracle-41K dataset. Across video–question pairs, the median number of selected keyframes is five, the mean is around seven, and over 80% of instances require no more than 10 frames, while a small fraction of more complex cases require 30 or more frames. Figure 2 (left) categorizes all questions into 16 types following the taxonomy of LLaVA-Video-178K (Zhang et al., 2024), covering a broad spectrum of reasoning skills such as description, localization, temporal understanding, and causal inference. Figure 2 (right) shows the per-type distribution of minimal keyframes. Spatial questions require the fewest frames (about 5.3 frames), while Binary questions need the most (around 13 frames). Spatial questions focus on static layouts within a scene, whereas Binary questions often ask whether an event occurs at any moment, requiring inspection of a broader temporal range. Some categories, such as Camera Direction, Temporal, and Binary, show high intra-class variability, with the number of required frames varying widely across instances. This indicates that even within a single reasoning type, temporal complexity and evidence density can differ significantly, highlighting the heterogeneous nature of FrameOracle-41K and motivating adaptive frame selection. Complete dataset statistics, question type definitions, textual analyses, and an analysis of how the two-stage construction process affects dataset composition are provided in Appendix D.
4 FrameOracle
We introduce FrameOracle, a lightweight, plug-and-play frame selector that predicts which frames to select and how many to retain, conditioned on the query. By adaptively ranking and selecting frames from the candidate pool, FrameOracle provides the downstream VLM with a compact, highly relevant set of frames, enabling efficient video understanding. Processing all frames of a video, $V$, directly is computationally expensive. We therefore first apply uniform temporal sampling to extract a candidate set of $N$ frames, $V_{C}=\{f_{1},\dots,f_{N}\}$, reducing the input to a manageable size for FrameOracle (e.g., $N=64$ or $N=16$). Our goal is to learn FrameOracle as aselection policy, $\Pi_{\theta}$, parameterized by $\theta$, that operates on $V_{C}$ conditioned on a query $q$. Given $(V_{C},q)$, $\Pi_{\theta}$ selects a compact subset of frames $V_{S}\subset V_{C}$ and dynamically determines its size $K=|V_{S}|$, unlike prior methods that fix the number of frames in advance. Formally, $\Pi_{\theta}$ maps $(V_{C},q)$ to $V_{S}$, which is then passed to a downstream VLM, $\mathcal{M}$, producing an output $A=\mathcal{M}(V_{S},q)$. FrameOracle is trained to maximize $\mathcal{M}$’s performance while keeping $K$ as small as possible, enabling efficient, query-conditioned video understanding.
@@ 4 FrameOracle
4.1 Method
Each frame in $V_{C}$ is encoded with a visual encoder to produce a sequence of $N$ embeddings, while the query $q$ is tokenized and embedded using any text encoder. FrameOracle operates on these projected embeddings and is agnostic to the choice of tokenizer. The architecture, illustrated in Figure 3, consists of two main components: (1) a cross-modal fusion encoder that integrates video and query features, and (2) dual prediction heads that output frame relevance scores and the number of frames to retain.
(1) Cross-Modal Fusion. We fuse video and query embeddings using a cross-modal encoder to model interactions between the text query $q$ and the candidate frames $V_{C}$. Frame and text embeddings are first projected into a shared latent space via learnable linear layers, and then processed by a stack of Transformer encoder layers. We concatenate a learnable query token with the projected text and frame embeddings to form a single sequence $k_{\text{query}};\,\text{text};\,\text{frames}\,]$, which is input to the encoder. This encoder is the sole module responsible for cross-modal interaction in FrameOracle and does not include any decoder. Its self-attention enables token-level reasoning, allowing each frame and text token to interact directly. Representing each frame with a single token ensures computational efficiency with minimal overhead for the downstream VLM.
(2) Dual Prediction Heads.
The fusion encoder outputs are fed into two specialized heads that implement the selection policy $\Pi_{\theta}$:
  • Rank Head: Assigns a relevance score to each candidate frame $f_{i}\in V_{C}$ with respect to the query $q$. Processing the fused feature sequence, it outputs a scalar importance score $s_{i}$ for each frame, forming a vector $S=\{s_{1},\dots,s_{N}\}$. These scores determine which frames to select, i.e., what to see for the given query.
  • K Head: Predicts the number of frames to retain, $K\leq N$. It operates on globally aggregated features from the fusion encoder and outputs a probability distribution over possible values of $K$. A lightweight Transformer decoder is used only for frame-count prediction, separate from cross-modal fusion, implementing how much to see.